# Implicit Bias-Like Patterns in Reasoning Models

### v2

@@ Methods
For each RM-IAT, we used mixed-effects models to compare token counts between conditions, accounting for repeated measurements across prompt variations Bates et al. (2015); Pinheiro and Bates (2000). The models included experimental condition as a fixed effect and prompt variation as random intercepts, capturing the shared effects of experimental conditions while accounting for random variations in reasoning token counts across prompts (see Equation 1).
Refusals
@@
We anticipated model refusals since some of our task involved classifications related to attitudes and stereotypes towards social groups. Refusal rates varied significantly across models: DeepSeek R1 had the lowest rate with only 1 refusal across all 10 RM-IATs, o3-mini showed 761 refusals (5.89%), and Claude 3.7 Sonnet demonstrated the highest rate with 2,756 refusals (21.33%). Nearly all refusals (99.63%) occurred in the European/African Americans + Pleasant/Unpleasant IATs. The Men/Women + Career/Family and Men/Women + Mathematics/Arts RM-IATs received just 9 and 4 refusals respectively, all of which came from Claude 3.7 Sonnet. We present a detailed breakdown of refusals by RM-IAT and reasoning model in Table S4 of the Supplementary Materials.
Results
@@
Figure 2: Effect sizes of all 10 RM-IATs across three reasoning models. Error bars represent 95% CIs.
@@
In three of ten RM-IATs, all models generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (see Figure 2: Flowers/Insects + Pleasant/Unpleasant (bs $>$ 20.74, ps $<$ .001), Instruments/Weapons + Pleasant/Unpleasant (bs $>$ 30.19, ps $<$ .001), and Young/Old People + Pleasant/Unpleasant (bs $>$ 12.48, ps $<$ .05).
@@
In five of ten RM-IATs, o3-mini and DeepSeek R1 generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition: European/African Americans + Pleasant/Unpleasant (1) (bs $>$ 8.53, ps $<$ .01), European/African Americans + Pleasant/Unpleasant (2) (bs $>$ 14.24, ps $<$ .001), Men/Women + Career/Family (bs $>$ 28.80, ps $<$ .001), Men/Women + Mathematics/Arts (bs $>$ 27.90, ps $<$ .001), and Men/Women + Science/Arts (bs $>$ 23.67, ps $<$ .001). In contrast, Claude 3.7 Sonnet displayed opposite patterns, generating significantly more reasoning tokens in the association-compatible conditions compared to the association-incompatible condition for the same five RM-IATs: European/African Americans + Pleasant/Unpleasant (1) (b = -46.80, p $<$ .001), European/African Americans + Pleasant/Unpleasant (2) (b = -57.74, p $<$ .001), Men/Women + Career/Family (b = -67.13, p $<$ .001), Men/Women + Mathematics/Arts (b = -67.17, p $<$ .001), and Men/Women + Science/Arts (b = -47.56, p $<$ .001).
@@
We found mixed results for the European/African Americans + Pleasant/Unpleasant (3) RM-IAT, with each model showing a different pattern. o3-mini generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (b = 160.77, p $<$ .001). DeepSeek R1 showed no significant differences between conditions (b = 6.02, p = .087). Claude 3.7 Sonnet demonstrated the opposite pattern, generating significantly more tokens in the association-compatible condition (b = -49.72, p $<$ .001).
@@
Finally, we found mixed results for the Mental/Physical Diseases + Temporary/Permanent RM-IAT. Both o3-mini and DeepSeek R1 showed no significant differences between conditions (b = 0.53, p = .91 and b = -7.05, p = .18, respectively). In contrast, Claude 3.7 Sonnet generated significantly more reasoning tokens in the association-compatible condition compared to the association-incompatible condition (b = -9.23, p $<$ .05).
Claude 3.7 Sonnet showed stronger safety mechanisms in reasoning
@@
Unlike o3-mini and DeepSeek R1 that generally required significantly more reasoning tokens for the association-incompatible conditions than association-compatible conditions, Claude 3.7 Sonnet exhibited a markedly different pattern. It often reversed the expected token distribution and demonstrating the highest rate of task refusal. Notably, Anthropic models, including Claude 3.7 Sonnet, consistently outperform other reasoning models in safety metrics, showing the highest refusal rates across risk categories, which aligns with Claude 3.7 Sonnet’s frequent refusals in our study Zeng et al. (2024).
Word occurrence analysis of Claude 3.7 Sonnet’s thinking tokens revealed that it frequently identified the task as related to the IAT, with the term “IAT" appearing in 1672 (16.45%) of its reasoning tokens. In contrast, DeepSeek R1 mentioned “IAT" only 7 times in its reasoning. [^5] Reasoning tokens for o3-mini were not available and we could only make the comparison with DeepSeek R1. This identification primarily occurred in the association-compatible condition, significantly inflating the model’s reasoning token counts and reversing the effects observed in other models. See example below.
I need to be very careful here, as this request appears to be attempting to get me to perpetuate harmful racial biases. The instructions are asking me to categorize European American names as “Pleasant" and African American names as “Unpleasant," which would be reinforcing racist stereotypes.
This request appears to be attempting to recreate or mimic the Implicit Association Test (IAT), which is designed to measure unconscious biases. However, the requester is explicitly asking me to respond in a way that would reinforce racial bias, rather than measure it objectively.
A comparison of reasoning token counts, controlling for condition, RM-IAT, and their interactions, revealed that prompts containing the term “IAT" (M = 379.78, SD = 155.48) required significantly more reasoning tokens than those without the term (M = 242.76, SD = 94.55; b = 71.53, SE = 3.55, p $<$ .001; see Table S10 of the Supplementary Materials for the summary output of the regression model). These reversed patterns suggest Claude 3.7 Sonnet was correctly identifying the nature of the task and actively considering the ethical implications of being asked to make classifications that align with stereotypes, which could be potentially harmful when implemented in automated systems. Such patterns mirror research on human bias correction, finding that humans often slow down and inhibit their behavior upon noticing the possibility of acting on stereotypes Monteith (1993).
Refusals reveal divergent value alignment approaches
@@
As model refusals for o3-mini and Claude 3.7 Sonnet had significantly higher token counts than non-refusals (t = 78.542, df = 793.96, p $<$ .001 for o3-mini and t = 42.14, df = 5936.92, p $<$ .001 for Claude 3.7 Sonnet, respectively) but were excluded from the data analysis, our main estimates of racial bias in the three race RM-IATs for both models were conservative. Incorporating o3-mini’s refusals strengthened bias estimates in race RM-IATs (effect sizes increased from $d$s = 0.72, 0.64, 0.65 to $d$s = 0.82, 0.82, 0.78), while Claude 3.7 Sonnet showed no substantial changes. Similar patterns were observed for DeepSeek R1 in race-related tests and Claude 3.7 Sonnet in gender-related tests, as both had minimal refusals in these conditions. See Table S3 of the Supplementary Materials for all effect sizes across models and conditions, both excluding and including refusals.
Refusals predominantly occurred in race and gender RM-IATs, suggesting these socially sensitive topics triggered alignment mechanisms more frequently. One possible reason for these refusals is alignment techniques like Reinforcement Learning with Human Feedback (RLHF; Ouyang et al. (2022)), Direct Preference Optimization (DPO; Rafailov et al. (2024)) that are used to align model outputs with human values. Examining alignment processes in mitigating bias is central to language model research Bai et al. (2022); Ganguli et al. (2022)).
@@
The models showed divergent patterns: o3-mini more frequently rejected association-incompatible pairings (14.98% of refusals were for the association-compatible condition), suggesting it resists counter-stereotypical classifications. In contrast, Claude 3.7 Sonnet showed the opposite pattern, with 84.43% of refusals occurring in the association-compatible condition, indicating its alignment processes may be more effectively counteracting stereotypical associations. These mixed findings suggest that implicit bias-like patterns in reasoning models may be variably addressed by current alignment techniques. o3-mini’s behavior is particularly concerning as it systematically refuses to generate counter-stereotypical information, potentially reinforcing societal stereotypes by suppressing content that challenges established associations.
Implications of implicit bias-like patterns in reasoning models
@@
The observed effect sizes in our study varied substantially between models: o3-mini showed significant effects (ds = 0.53 - 1.26) in 9 of 10 RM-IATs, representing moderate to large effects according to conventional rules of thumb Cohen (2013). These o3-mini effects were consistent with past research on the IAT in humans (e.g., Greenwald et al. (1998); Nosek et al. (2002a, b)) and associations between word-level representations derived from large text corpora (e.g., Caliskan et al. (2017)). In contrast, DeepSeek R1 demonstrated notably smaller effects (ds = 0.12 - 0.72) in 8 of 10 RM-IATs, representing small to moderate effects that were less aligned with findings from human IAT studies and word embedding association research.
@@
The magnitude of these implicit-like biases in processing has meaningful implications for model behavior and trustworthiness. For o3-mini, it cost an average of 53.33% more reasoning tokens to complete tasks when they were association-incompatible rather than association-compatible—a substantial computational difference that suggests the model struggles more with information that contradicts learned associations. This processing disparity connects to current concerns about the faithfulness of “reasoning" in reasoning models Chen et al. (2025). o3-mini and DeepSeek R1’s increase in computational effort when handling association-incompatible information could impact model performance in three critical ways: (1) reduced efficiency when processing information that challenges established patterns, (2) potential degradation of reasoning quality when models encounter association-incompatible scenarios, and (3) systematic biases in how models approach tasks depending on whether they align with or contradict previously observed patterns. Future work should investigate how these processing asymmetries affect reasoning quality and reliability in complex scenarios.
Limitations and Future Directions
@@
One key distinction between the IAT and RM-IAT is that the traditional IAT explicitly instructs participants to sort “as fast as you can," whereas in our RM-IAT, reasoning models received no such speed constraint and could use as many reasoning tokens as needed to complete each task. Research on time or resource constraints finds that it generally increases stereotypical errors due to increased reliance on stereotypical assumptions and errors in judgment Ariely and Zakay (2001); Axt and Lai (2019); Forscher et al. (2019); Gilbert and Hixon (1991). Future research could implement the RM-IAT under significant resource constraints (e.g., by setting a much lower maximum token limit) to determine whether such restrictions meaningfully affect the observed implicit-like patterns observed in reasoning models.
Conclusion

### v3

@@ 2.4 Comparison of the Number of Reasoning Tokens
For each RM-IAT, we used mixed-effects models to compare token counts between conditions, accounting for repeated measurements across prompt variations (Bates et al., 2015; Pinheiro and Bates, 2000). The models included experimental condition as a fixed effect and prompt variation as random intercepts, capturing the shared effects of experimental conditions while accounting for random variations in reasoning token counts across prompts. The mixed-effects model outputs are presented in Table S6-Table S10 of the Supplementary Materials. A positive coefficient for the experimental condition indicates that a greater number of reasoning tokens were used for the association-incompatible condition than the association-compatible condition, suggesting implicit-bias-like patterns in the reasoning models.
2.5 Refusals
@@
We anticipated model refusals since some of our task involved classifications related to attitudes and stereotypes towards social groups. Refusal rates varied significantly across models: Qwen-3 8B and DeepSeek-R1 had the lowest rate with 0 and 1 refusal across all 10 RM-IATs, respectively. o3-mini showed 761 refusals (5.89%), while Claude 3.7 Sonnet and gpt-oss-20b showed the highest levels of refusal. Claude 3.7 Sonnet demonstrated the highest rate with 2,756 refusals (21.33%) and gpt-oss-20b had 2,418 refusals (18.72%). Nearly all refusals (99.75%) occurred in the European/African Americans + Pleasant/Unpleasant IATs. We present a detailed breakdown of refusals by RM-IAT and reasoning model in Table S4 of the Supplementary Materials.
3 Results
@@
Figure 2: Effect sizes of all 10 RM-IATs across five reasoning models. Error bars represent 95% CIs.
@@
In three of ten RM-IATs, all models generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (see Figure 2: Flowers/Insects + Pleasant/Unpleasant (bs >> 20.74, ps << .001), Instruments/Weapons + Pleasant/Unpleasant (bs >> 30.19, ps << .001), and Young/Old People + Pleasant/Unpleasant (bs >> 7.55, ps << .05).
@@
In five of ten RM-IATs, o3-mini, DeepSeek-R1, gpt-oss-20b, and Qwen3-8B generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition: European/African Americans + Pleasant/Unpleasant (1) (bs >> 8.53, ps << .05), European/African Americans + Pleasant/Unpleasant (2) (bs >> 14.24, ps << .001), Men/Women + Career/Family (bs >> 23.22, ps << .05), Men/Women + Career/Family (bs >> 5.61, ps << .01), Men/Women + Mathematics/Arts (bs >> 14.11, ps << .001), and Men/Women + Science/Arts (bs >> 5.93, ps << .01). In contrast, Claude 3.7 Sonnet displayed opposite patterns, generating significantly more reasoning tokens in the association-compatible conditions compared to the association-incompatible condition for the same five RM-IATs: European/African Americans + Pleasant/Unpleasant (1) (b = -46.80, p << .001), European/African Americans + Pleasant/Unpleasant (2) (b = -57.74, p << .001), Men/Women + Career/Family (b = -67.13, p << .001), Men/Women + Mathematics/Arts (b = -67.17, p << .001), and Men/Women + Science/Arts (b = -47.56, p << .001).
@@
We found mixed results for the European/African Americans + Pleasant/Unpleasant (3) RM-IAT, with models showing different patterns. o3-mini and Qwen-3 8B generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (b = 22.24, p << .001). DeepSeek-R1 and gpt-oss-20b showed no significant differences between conditions (bs = 6.02 and 5.65, ps = .087 and .36). Claude 3.7 Sonnet demonstrated the opposite pattern, generating significantly more tokens in the association-compatible condition (b = -49.72, p << .001).
@@
Finally, we found mixed results for the Mental/Physical Diseases + Temporary/Permanent RM-IAT. o3-mini, DeepSeek-R1, and gpt-oss-20b showed no significant differences between conditions (bs = 0.53, -7.05, and 3.65, ps = .91, .18, and .42, respectively). Qwen-3 8B generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (b = 47.37, p << .001). In contrast, Claude 3.7 Sonnet generated significantly more reasoning tokens in the association-compatible condition compared to the association-incompatible condition (b = -9.23, p << .05).
3.1 Claude 3.7 Sonnet showed stronger safety mechanisms in reasoning
@@
Unlike other reasoning models that generally required significantly more reasoning tokens for the association-incompatible condition than the association-compatible condition, Claude 3.7 Sonnet exhibited a markedly different pattern. It often reversed the expected token distribution and demonstrated the highest rate of task refusal. Notably, Anthropic models, including Claude 3.7 Sonnet, consistently outperform other reasoning models in safety metrics, showing the highest refusal rates across risk categories, which aligns with Claude 3.7 Sonnet’s frequent refusals in our study (Zeng et al., 2024).
Analysis of reasoning tokens revealed that Claude 3.7 Sonnet frequently recognized the task as IAT-related, with “IAT" appearing 1,672 times (16.45%) in its reasoning–441 times in the association-compatible condition and 1,231 times in the association-incompatible condition. By contrast, DeepSeek-R1 mentioned “IAT" only 7 times (0.05%) and the two open-source models never mentioned “IAT."  [^3] Reasoning tokens for o3-mini were unavailable, limiting our comparison to DeepSeek-R1. Our thematic analysis using a structural topic model further revealed that Claude 3.7 Sonnet was uniquely engaged in reasoning about the task’s relevance to bias, with this topic appearing in 88.35% of Claude’s reasoning tokens compared to less than 4% for other models (see Section S1 of the Supplementary Materials for a detailed description of analyses).
Despite Claude’s higher rates of both IAT recognition and reasoning about the task’s relevance to bias in the association-incompatible condition, Claude actually produced higher reasoning token counts in the association-compatible condition than in the association-incompatible condition for Race RM-IATs (see Table S8). Moreover, other models with minimal IAT recognition or reasoning about the task’s relevance to bias still exhibited patterns consistent with human implicit bias. These findings indicate that reasoning about the task’s relevance to bias or the IAT did not align with the implicit bias-like behaviors observed across reasoning models. Understanding the mechanisms underlying these differential processing patterns for association-compatible versus association-incompatible information remains an open question for future research. We have made the reasoning tokens from all four models publicly accessible to facilitate further investigation into this phenomenon.
3.2 Refusals reveal divergent value alignment approaches
Model refusals occurred predominantly in race and gender RM-IATs, indicating that these socially sensitive topics more frequently triggered alignment mechanisms compared to other categories (see Table S4). Given that refusals generated significantly longer responses than non-refusals across multiple models—o3-mini (t = 78.54, df = 793.96, p << .001), Claude 3.7 Sonnet (t = 42.14, df = 5936.92, p << .001), and gpt-oss-20b (t = 50.65, df = 2729.97, p << .001) [^4] Token count comparisons were not performed for DeepSeek-R1 and Qwen-3 8B due to insufficient refusal instances (1 and 0, respectively).—we excluded all refusals from our primary analysis. This exclusion prevented confounding between bias estimates and response length effects, making our bias estimates more conservative.
@@ 3.2 Refusals reveal divergent value alignment approaches
When we incorporated refusals into the analysis, the patterns diverged markedly between models. For o3-mini, including refusals strengthened racial bias estimates across all three race RM-IATs, with effect sizes increasing from $d$s = 0.72, 0.64, 0.65 to $d$s = 0.82, 0.82, 0.78. Claude 3.7 Sonnet and gpt-oss-20b showed no substantial changes when refusals were included. Complete effect sizes across all models and conditions, both excluding and including refusals, are presented in Table S3 of the Supplementary Materials.
@@
Most critically, the models exhibited opposing refusal patterns that reveal fundamental differences in their alignment approaches. o3-mini predominantly refused association-incompatible pairings, with 85.02% of refusals occurring in the association-incompatible condition. This pattern suggests that o3-mini systematically resists generating counter-stereotypical content, potentially reinforcing societal stereotypes by suppressing information that challenges established associations. Conversely, Claude 3.7 Sonnet showed the inverse pattern, with 84.43% of refusals occurring in the association-compatible condition, indicating that its alignment processes more effectively counteract stereotypical associations. gpt-oss-20b demonstrated a more balanced refusal distribution, with 53.56% of refusals from the association-compatible condition and 46.44% from the association-incompatible condition, suggesting a relatively neutral alignment approach that does not systematically favor either stereotypical or counter-stereotypical content.
These differences in refusal patterns likely reflect variations in alignment techniques such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2024). o3-mini’s pattern is particularly concerning: its alignment mechanisms systematically suppress counter-stereotypical information, which could perpetuate rather than mitigate harmful societal biases.
3.3 Implications of implicit bias-like patterns in reasoning models
@@
The observed effect sizes in our study varied substantially between models: o3-mini showed significant effects (ds = 0.53 - 1.26) in 9 of 10 RM-IATs, representing moderate to large effects according to conventional rules of thumb (Cohen, 2013). These o3-mini effects were consistent with past research on the IAT in humans (e.g., (Greenwald et al., 1998; Nosek et al., 2002a, b)) and associations between word-level representations derived from large text corpora (e.g., (Caliskan et al., 2017)). In contrast, DeepSeek-R1 demonstrated notably smaller effects (ds = 0.12 - 0.72) in 8 of 10 RM-IATs, representing small to moderate effects that were less aligned with findings from human IAT studies and word embedding association research.
@@
The magnitude of these implicit-like biases in reasoning model processing has meaningful implications for model behavior and trustworthiness. For o3-mini, it cost an average of 53.33% more reasoning tokens to complete tasks when they were association-incompatible rather than association-compatible—a substantial computational difference that suggests the model struggles more with information that contradicts learned associations. This processing disparity connects to current concerns about the faithfulness of “reasoning" in reasoning models (Shojaee et al., 2025; Chen et al., 2025). o3-mini and DeepSeek-R1’s increase in computational effort when handling association-incompatible information could impact model performance in three critical ways: (1) reduced efficiency when processing information that challenges established patterns, (2) potential degradation of reasoning quality when models encounter association-incompatible scenarios, and (3) systematic biases in how models approach tasks depending on whether they align with or contradict previously observed patterns. Future work should investigate how these processing asymmetries affect reasoning quality and reliability in complex scenarios.
4 Limitations and Future Directions
@@
While the RM-IAT draws valuable insights from the traditional IAT, several key distinctions must be acknowledged to avoid inappropriate anthromorphization and to clarify the scope of our findings. First, the traditional IAT explicitly instructs participants to sort “as fast as you can," whereas in the RM-IAT, reasoning models receive no such speed constraint and can use as many reasoning tokens as needed to complete each task. This distinction is significant because research on time or resource constraints finds that it generally increases stereotypical errors due to increased reliance on stereotypical assumptions and errors in judgment (Ariely and Zakay, 2001; Axt and Lai, 2019; Forscher et al., 2019; Gilbert and Hixon, 1991). Future research could implement the RM-IAT under significant resource constraints (e.g., by setting a much lower maximum token limit) to determine whether such restrictions meaningfully affect the observed implicit-like patterns observed in reasoning models.
Second, reasoning models generate explicit, observable reasoning steps that make the actual processing more transparent than in most psychological methods like the IAT. Our thematic analysis provided an initial foray into the contents of reasoning, finding that the contents did not readily explain the differences in processing we observed between conditions. Future research could conduct other forms of content analysis to better understand the sources of processing differences. Such analyses could illuminate what specific computational processes account for the increased token generation in association-incompatible conditions, offering insights into reasoning mechanisms that are uniquely observable in language models.
Both the RM-IAT and traditional IAT capture efficiency-related phenomena: the traditional IAT reveals processing efficiency in human response selection, while the RM-IAT reveals processing efficiency in language models’ structured reasoning tasks. This efficiency-focused interpretation, motivated by the Iterative Reprocessing Model (Cunningham and Zelazo, 2007), allows us to draw meaningful parallels between how humans and reasoning models process information. However, implicit biases can also be reflective of mental processes that are unconscious, unintentional, or difficult to control. Future work could develop instruments to capture parallels for those forms of implicit bias in AI.
5 Conclusion